OpenStudy (anonymous):

a room mesures 12 1/4 by10 10 1/2 hat is the area of the room

OpenStudy (anonymous):

the measurements dont sound right. post your question clearly

OpenStudy (anonymous):

a room measure wait sry i dnt know howtoput the farctions bar under

OpenStudy (anonymous):

what does 10 101/2 mean?

OpenStudy (anonymous):

oh ok correction 10 1/2

OpenStudy (anonymous):

okay. so what is the area of a rectangle?

OpenStudy (anonymous):

yes thats the question to my problem

OpenStudy (anonymous):

right, so do you know the formula to find the area of a rectangle of length L and width W?

OpenStudy (anonymous):

yeah leath times width jstnt know how tosle usingmxed numbers

OpenStudy (anonymous):

12 1/4 means twelve and a quarter. it means 12 +1/4 = 12+0.25 = 12.25

OpenStudy (anonymous):

that not wait you are confusing me kow

OpenStudy (anonymous):

what is the confusion about? 12 1/4 just means twelve and one quarter. 2 1/2 means two and a half. so 2 1/2 means 2.5

OpenStudy (anonymous):

well because my answer chooses are 120 3/4 124 1/2 127 1/2 128 5/8

OpenStudy (anonymous):

do you know how to converted mixed fractions to normal fractions?

OpenStudy (anonymous):

http://www.youtube.com/watch?v=aL3znpsf8X0